Cummins India Ltd has the following capital structure, which it considers optimal:

Debt 25%

Preference Shares 10%

Equity shares 65%

Total 100%

Applicable tax rate for the company is 25%. Risk free rate of return is 6%, average equity

market investment has expected rate of return of 12%. The company’s beta is 1.10.

Following terms would apply to new securities being issued as follows:

1. New preference can be issued at a face value of Rs. 100 per share, dividend and cost of

issuance will be Rs. 10 per share and Rs. 2 per share respectively.

2. Debt will bear an interest rate of 9%.

Calculate

a. component cost of debt, preference shares and equity shares assuming that the company

does not issue any additional equity shares.

 

b. WACC.


1
Expert's answer
2021-05-26T13:41:17-0400

a. Component cost of debt is:

"0.1\u00d7(1 - 0.25) = 0.075" or 7.5%.

Cost of preference shares is:

"12\/100 = 0.12" or 12%.

and cost of equity shares is:

"Ke = 0.06 + 1.5\u00d7(0.15 - 0.06) = 0.195" or 11.25%.

b. "WACC = 0.195\u00d70.6 + 0.075\u00d70.25 + 0.12\u00d70.15 = 0.1538."